10. (10 points) A ball of mass 135 g is placed in contact with a horizontal spring, which is compressed by a distance of 12 cm. The spring constant of this particular spring is 384 N/m (a) Evaluate the elastic potential energy stored in the compressed spring. (b) Now the spring is released, launching the ball horizontal. Find the speed of the ball immediately after it is launched.
Answer #1

a)
compression = 12 cm
= 0.12 m

use,
potential energy = 0.5*k*(compression)^2
= 0.5*384*(0.12)^2
= 2.76 J

b)
when ball will be launched
potential energy of spring will get transmited as kinetic energy of ball
so,
0.5*m*v^2 = potential energy
0.5*0.135*v^2 = 2.76
v^2 = 40.9
v = 6.39 m/s